# Ira: Efficient Transaction Replay for Distributed Systems

### v1

@@ 5.3.1 Per-Block Speedup Distribution
Figure˜13 shows per-block speedup across the entire block range for the sequential backup configuration; Table˜10 in the appendix lists the key percentiles for all three configurations.
Figure 13: Per-block speedup (sequential backup vs. baseline) across 100,800100,800{} blocks. Each point is one block; the solid line shows a $500$-block rolling average and the shaded band spans the P5–P95 range. The dashed line marks the median speedup.
The backup achieves a median speedup of $24.9{}\times$ (mean $28.9{}\times$) across the sequential configuration, with the parallel configurations yielding nearly identical distributions. In terms of aggregate wall time, the sequential backup completes all 100,800100,800{} blocks in 4,3664,366{} s ($5.2{}\times$ faster than the baseline’s 22,60422,604{} s), parallel-$16$ in 956956{} s ($23.6{}\times$), and parallel-$64$ in 871871{} s ($25.9{}\times$).
The speedup magnitude follows directly from Amdahl’s law and the I/O-dominance of the baseline. Using backup execution time as a proxy for pure compute (the backup’s cache is warm from prefetching), the baseline spends approximately $96\%$ of its time on page faults and only ${\sim}4\%$ on EVM computation. Eliminating all I/O yields a theoretical ceiling of $1/(1-0.96)=25\times$ which is almost exactly the observed median. Hint-driven prefetching approaches this ceiling because it resolves every storage key before execution begins, converting the baseline’s serial page faults into a single overlapped prefetch phase.
Figure˜13 reveals three regimes. First, an initial transient in the first ${\sim}500$ blocks where speedup exceeds $50\times$: the baseline starts with a fully cold OS page cache and mean execution time of $570$ ms, while the backup begins against an already-populated MDBX file and has zero wait time, completing in under $9$ ms. As the baseline’s page cache fills (RSS grows from $0.2$ GB to ${\sim}8$ GB over this window), its execution time drops and the speedup converges to steady state.
@@ 5.3.2 Tail Behavior
With sequential prefetching, $3.02{}\%$ of blocks experience a regression (Table˜11 in the appendix). The root cause is MDBX’s use of memory-mapped I/O: each cold page access blocks the faulting thread until the kernel completes the read (Section˜A.3), so a single prefetch thread issues at most one storage I/O at a time. When a block touches thousands of uncached B-tree pages, this serialization causes the executor to stall waiting for prefetch to finish, negating the I/O savings. Among these regressed blocks, the median speedup is $0.18\times$ ($5.5\times$ slower than baseline), with a mean of $0.20\times$ ($4.9\times$ slower). At $64$ prefetch threads, only $0.09{}\%$ of blocks remain slower than the baseline.
@@ 5.3.3 Wait Time Analysis
Figure˜14 decomposes aggregate wall time by whether the executor stalled waiting for prefetch to complete.
@@ 5.4 Thread Scaling
Figure˜15 shows the wall-time speedup relative to the baseline as a function of prefetch thread count across ten configurations (116464 threads). The curve exhibits three regimes. First, a steep scaling region from $1$ to $8$ threads where wall time drops from 4,3664,366{} s to ${\sim}1{,}088$ s ($4\times$ reduction), reflecting near-linear conversion of serial page faults into parallel I/O. Second, a diminishing-returns region from $8$ to $16$ threads where wall time falls further to 956956{} s ($4.6{}\times$ over sequential) as the SSD’s command queue approaches saturation. Third, a plateau beyond $16$ threads: doubling from $16$ to $32$ yields only a $7\%$ reduction, and quadrupling to $64$ threads reaches 871871{} s ($1.1{}\times$ over $16$ threads), indicating that the I/O bandwidth of our NVMe SSD is fully saturated.
@@ 5.5 Resource Usage
All three configurations consume comparable memory (${\sim}26.3{}$ GB), dominated by the MDBX memory-mapped database. The primary’s hint data structures (the next-access map) add negligible overhead: the map contains at most one entry per unique key in the current block (median 1,8061,806{} keys), each requiring a few tens of bytes, totaling at most a few hundred kilobytes per block.
@@
6 Discussion
@@
Deployment considerations. Layer-1 optional. Existing database systems can deploy Ira without enforcing hint correctness as part of the underlying synchronization protocol. Blockchains that implement this need not enforce correctness as part of the block proposal mechanism as the hints are advisory and affect only performance.
@@
7 Related Work
Optimal Cache Replacement. Belady’s MIN algorithm [beladyStudyReplacementAlgorithms1966], proved optimal by Mattson et al. [mattsonEvaluationTechniquesStorage1970], provides optimal offline cache replacement by evicting the item whose next access is furthest in the future. Some works use adaptive heuristics, e.g., the adaptive replacement cache [megiddoARCSelfTuningLow2003] (ARC) and LIRS [jiangLIRSEfficientLow2002] which use heuristics to dynamically balance between recency and frequency, and LeCaR [vietriDrivingCacheReplacement2018] uses regret minimization to learn this balance online. Another approach is to use hardware prefetchers which predict future accesses to hide memory latency. Prefetch-aware replacement policies like PACMan [wuPACManPrefetchawareCache2011] and SHiP [wuSHiPSignaturebasedHit2011] adjust eviction priorities for prefetched blocks. A different approach is to approximate the MIN algorithm from history. Works such as the HawkEye cache by Jain and Lin [jainRethinkingBeladysAlgorithm2018] observe that MIN minimizes total misses but not demand misses in the presence of prefetching. They introduce Demand-MIN and Flex-MIN to navigate this tradeoff. The Hawkeye cache [jainBackFutureLeveraging2016] uses past access patterns to approximate MIN decisions. Subsequent works like Mockingjay [shahEffectiveMimicryBeladys2022] and Glider [shiApplyingDeepLearning2019] use machine learning to improve prediction accuracy but require specialized hardware structures and still cannot exceed the accuracy ceiling imposed by predicting from past behavior alone. Recent deep learning approaches [shiHierarchicalNeuralModel2021, yangSGDPStreamGraphNeural2023, ganfureDeepPrefetcherDeepLearning2020] use neural networks to learn access patterns, achieving higher accuracy than traditional heuristics but remaining fundamentally limited to inference from history. Our work differs from these works in that we transmit exact future access information from primary to backup, enabling true optimal eviction without prediction.
Blockchain State Synchronization. Blockchain nodes face significant replay challenges during both initial synchronization and ongoing block validation. Early Ethereum clients performed full sync by re-executing every transaction from genesis, but this approach became impractical as chain history grew. Geth [EthereumGoethereum2026] introduced fast sync [Eth63Fast] (2015) and later snap sync [GethV1100] (2021), which download recent state directly and reconstruct the Merkle trie locally, achieving $10\times$ faster initial sync. Erigon [ErigontechErigon2026] (formerly Turbo-Geth) pioneered staged sync [ErigonEthStagedsync], a pipelined architecture that processes headers, bodies, senders recovery, and execution as discrete stages; Reth [ParadigmxyzReth2026] adopts this architecture in Rust for further performance gains. Despite these optimizations, the execution stage, where transactions are replayed to compute state, remains the primary bottleneck and cannot be parallelized within staged sync. To enable parallel validation, some systems require transactions to declare their data dependencies upfront. Ethereum’s EIP-2930 [proposalsEIP2930OptionalAccess] introduced optional access lists that pre-declare storage slots, though adoption remains low: analysis shows only $1.46\%$ of transactions include access lists despite $42.6\%$ benefiting from them [heimbachDissectingEIP2930Optional2023]. Solana’s Sealevel runtime [yakovenkoSolanaNewArchitecture2025] mandates read/write set declarations, enabling lock-based parallel execution across non-conflicting transactions. For systems without mandatory access declarations, Block-STM [gelashviliBlockSTMScalingBlockchain2022] provides optimistic parallel execution using software transactional memory techniques, achieving 170k transactions per second on Aptos. Sui [blackshearSuiLutrisBlockchain2024] takes an object-centric approach where owned objects bypass consensus entirely, enabling sub-second finality for non-shared state. Forerunner [chenForerunnerConstraintbasedSpeculative2021] exploits the time window between transaction dissemination and consensus to speculatively pre-execute transactions, achieving $10\times$ speedup through constraint-based multi-trace specialization. Looking forward, Verkle trees [kuszmaulVerkleTrees2018] promise to enable stateless clients by reducing witness sizes from ${\sim}150$ KB to 121-2 KB, allowing validators to verify blocks without storing full state. While these works optimize sync architecture, transaction scheduling, or state representation, none address cache eviction during replay. Our system Ira is the first complementary approach that provides exact future access information, enabling Belady-optimal caching during the execution stage improving upon the bottleneck that staged sync, access lists, and parallel execution engines aimed to solve.
Speculative and parallel execution. Parallel transaction execution has been extensively studied in both database and blockchain contexts. Kung and Robinson’s seminal work on optimistic concurrency control (OCC) [kungOptimisticMethodsConcurrency1981] established the read-validate-write paradigm that underlies modern parallel execution systems. Silo [tuSpeedyTransactionsMulticore2013] demonstrated that OCC can achieve nearly $700,000$ transactions per second on multicore hardware by eliminating centralized contention points. For workloads with high contention on popular records, Doppel [narulaPhaseReconciliationContended2014] introduced phase reconciliation, cycling between split phases (where commutative operations execute on per-core state) and joined phases (where results are merged), achieving up to $38\times$ higher throughput than conventional protocols. Deterministic approaches like BOHM [faleiroRethinkingSerializableMultiversion2015] pre-compute transaction dependencies to enable parallel execution without runtime conflict detection, while PWV [faleiroHighPerformanceTransactions2017] increases concurrency by making writes visible before transaction commit through piece-wise execution. In the blockchain space, the previous paragraph discussed systems like Block-STM, Forerunner, and access-list-based parallelization; commercial efforts such as Monad [IntroductionMonadDeveloper] and Sei [SeiParallelizationEngine2026] further pursue optimistic parallel EVM execution. These works share a common goal: maximizing transaction throughput by exploiting parallelism. Ira is orthogonal to and composable with all of these approaches. Even when transactions execute in parallel, each execution thread must still manage its cache and make eviction decisions. By providing exact future access information, Ira enables Belady-optimal caching within each parallel execution context, reducing I/O latency regardless of the concurrency control protocol employed.
8 Conclusion
Appendix A Reth Architecture
@@
Ethereum uses two clients: a consensus client that follows a beacon chain to determine the blockchain that is valid, and an execution client that executes the blocks from the consensus client. They work together to keep a node in sync, validate incoming blocks, and to propose new blocks. This section describes the architecture of Reth, a Rust-based high performance execution client, in particular, we focus on the block validation and execution architecture relevant to understanding our hint-based optimization. Figure˜16 presents a schematic summary of the architecture that we are interested in.
A.1 System components
@@ A.3 MDBX database
A.4 Reth Baseline Architecture
@@ Appendix B Ira-L Architecture
B.1 Reth Primary
reth-primary is the hint generator that simulates a block proposer. It replays blocks, collects all state keys accessed during execution, and writes hints that enable other backups to prefetch state for faster replay. Figure˜17 presents an illustration of our architecture. Components.
@@ B.1 Reth Primary
B.2 Reth Backup
reth-backup is the hint consumer that simulates a validator receiving hints from the proposer. It prefetches all required state before execution, so that block processing incurs zero disk I/O. Figure˜18 presents an illustration of our architecture.
@@ B.2 Reth Backup
Appendix C Additional Details
Table˜5 provides detailed bucketed statistics for the intra-block access frequency distribution shown in Figure˜5. Table˜6 provides the detailed breakdown for the contract concentration shown in Figure˜6. Table˜7 provides the detailed breakdown for the key lifespan distribution shown in Figure˜7. Table˜8 provides per-block statistics shown in Figure˜8. Section˜C.1 provides the source annotation distribution referenced in Section˜4.1. Section˜C.2 provides the tabulated backup speedup and tail behavior data. Section˜C.3 provides a detailed analysis of hint cost scaling and variance.
@@ Appendix C Additional Details
C.1 Source Annotation Distribution
@@
C.2 Backup Speedup Details
Table˜10 provides the per-block speedup percentiles summarizing Figure˜13. Table˜11 provides the tail behavior breakdown referenced in Section˜5.3.2.
@@ C.2 Backup Speedup Details
C.3 Hint Cost Variance Analysis
The per-block hint fraction (Figure˜9) varies across blocks despite a stable per-key cost. Table˜12 shows how the hint fraction changes with block working set size. The median hint fraction decreases monotonically from $24.8{}\%$ for small blocks (<<500 keys) to $9.7{}\%$ for large blocks (>>5K keys), even though absolute hint time increases. This occurs because execution time grows faster than hint time for larger blocks: large blocks perform more I/O, and the I/O-dominated execution time scales super-linearly with working set size while hint construction scales linearly (Figure˜10).

### v2

@@ 5.3.1 Per-Block Speedup Distribution
Figure˜11 shows per-block speedup across the entire block range for the sequential backup configuration; Table˜10 in the appendix lists the key percentiles for all three configurations.
Figure 11: Per-block speedup (sequential backup vs. baseline) over the 100,800100,800{}-block trace. Each point is one block; the solid line is a $500$-block rolling mean and the shaded band spans the rolling P5–P95 within each window. The dashed line marks the overall median ($24.9{}\times$).
The backup achieves a median per-block speedup of $24.9{}\times$ (mean $28.9{}\times$) across the sequential configuration, with the parallel configurations yielding nearly identical distributions. In terms of aggregate wall time, the sequential backup completes all 100,800100,800{} blocks in 4,3664,366{} s ($5.2{}\times$ faster than the baseline’s 22,60422,604{} s), parallel-$16$ in 956956{} s ($23.6{}\times$), and parallel-$64$ in 871871{} s ($25.9{}\times$). The gap between per-block and wall-time speedup for the sequential configuration arises because the $3.02{}\%$ of blocks with regressions (Section˜5.3.2) contribute disproportionately to aggregate wall time.
The speedup magnitude reflects the I/O-dominance of baseline execution (Section˜3): because storage I/O accounts for the majority of block execution time, prefetching the entire working set before execution begins eliminates nearly all of it, leaving only EVM computation.
Figure˜11 reveals three regimes. First, an initial transient in the first ${\sim}500$ blocks where speedup exceeds $50\times$: the baseline starts with a fully cold OS page cache, while the backup begins against an already-populated MDBX file and completes with zero wait time. As the baseline’s page cache fills, its execution time drops and the speedup converges to steady state.
@@ 5.3.2 Tail Behavior
With sequential prefetching, $3.02{}\%$ of blocks experience a regression (Table˜11 in the appendix). The root cause is MDBX’s use of memory-mapped I/O: each cold page access blocks the faulting thread until the kernel completes the read (Section˜C.3), so a single prefetch thread issues at most one storage I/O at a time. When a block touches thousands of uncached B-tree pages, this serialization causes the executor to stall waiting for prefetch to finish, negating the I/O savings. Among these regressed blocks, the median speedup is $0.18\times$ ($5.5\times$ slower than baseline), with a mean of $0.20\times$ ($4.9\times$ slower). At $64$ prefetch threads, only $0.09{}\%$ of blocks remain slower than the baseline.
@@ 5.3.3 Wait Time Analysis
Figure˜12 decomposes aggregate wall time by whether the executor stalled waiting for prefetch to complete.
@@ 5.4 Thread Scaling
Figure˜13 shows the wall-time speedup relative to the baseline as a function of prefetch thread count across ten configurations (116464 threads). The curve exhibits three regimes. First, a steep scaling region from $1$ to $8$ threads where wall time drops from 4,3664,366{} s to ${\sim}1{,}088$ s ($4\times$ reduction), reflecting near-linear conversion of serial page faults into parallel I/O. Second, a diminishing-returns region from $8$ to $16$ threads where wall time falls further to 956956{} s ($4.6{}\times$ over sequential) as the SSD’s command queue approaches saturation. Third, a plateau beyond $16$ threads: doubling from $16$ to $32$ yields only a $7\%$ reduction, and quadrupling to $64$ threads reaches 871871{} s ($1.1{}\times$ over $16$ threads), indicating that the I/O bandwidth of our NVMe SSD is fully saturated.
@@ 5.5 Resource Usage
All three configurations consume comparable memory (${\sim}26.3{}$ GB), dominated by the MDBX memory-mapped database. The primary’s hint data structures (the access set) add negligible overhead: the set contains at most one entry per unique key in the current block (median 1,8061,806{} keys), each requiring a few tens of bytes, totaling at most a few hundred kilobytes per block.
@@
6 Related Work
Optimal Cache Replacement. Belady’s MIN algorithm [beladyStudyReplacementAlgorithms1966], proved optimal by Mattson et al. [mattsonEvaluationTechniquesStorage1970], provides optimal offline cache replacement by evicting the item whose next access is furthest in the future. Some works use adaptive heuristics, e.g., the adaptive replacement cache [megiddoARCSelfTuningLow2003] (ARC) and LIRS [jiangLIRSEfficientLow2002] which use heuristics to dynamically balance between recency and frequency, and LeCaR [vietriDrivingCacheReplacement2018] uses regret minimization to learn this balance online. Another approach is to use hardware prefetchers which predict future accesses to hide memory latency. Prefetch-aware replacement policies like PACMan [wuPACManPrefetchawareCache2011] and SHiP [wuSHiPSignaturebasedHit2011] adjust eviction priorities for prefetched blocks. A different approach is to approximate the MIN algorithm from history. The Hawkeye cache [jainBackFutureLeveraging2016] uses past access patterns to approximate MIN decisions. Jain and Lin [jainRethinkingBeladysAlgorithm2018] later observe that MIN minimizes total misses but not demand misses in the presence of prefetching, introducing Demand-MIN and Flex-MIN to navigate this tradeoff. Subsequent works like Mockingjay [shahEffectiveMimicryBeladys2022] and Glider [shiApplyingDeepLearning2019] use machine learning to improve prediction accuracy but require specialized hardware structures and still cannot exceed the accuracy ceiling imposed by predicting from past behavior alone. Recent deep learning approaches [shiHierarchicalNeuralModel2021, yangSGDPStreamGraphNeural2023, ganfureDeepPrefetcherDeepLearning2020] use neural networks to learn access patterns, achieving higher accuracy than traditional heuristics but remaining fundamentally limited to inference from history. Our work differs from these works in that we transmit exact future access information from primary to backup, enabling true optimal eviction without prediction.
Blockchain State Synchronization. Blockchain nodes face significant replay challenges during both initial synchronization and ongoing block validation. Modern Ethereum clients use staged sync architectures [ErigonEthStagedsync, ParadigmxyzReth2026] that pipeline header, body, and execution stages; despite these optimizations, the execution stage remains the primary bottleneck. To enable parallel validation, some systems require transactions to declare data dependencies upfront: Ethereum’s EIP-2930 [proposalsEIP2930OptionalAccess] introduced optional access lists, though adoption remains low [heimbachDissectingEIP2930Optional2023]; Solana’s Sealevel runtime [yakovenkoSolanaNewArchitecture2025] mandates read/write set declarations. For systems without mandatory declarations, Block-STM [gelashviliBlockSTMScalingBlockchain2022] provides optimistic parallel execution, and Forerunner [chenForerunnerConstraintbasedSpeculative2021] exploits speculative pre-execution. While these works optimize sync architecture or transaction scheduling, none address cache eviction during replay. Ira provides exact future access information, enabling optimal caching during the execution stage. Appendix˜F provides extended discussion.
Speculative and parallel execution. Parallel transaction execution has been extensively studied. Kung and Robinson’s optimistic concurrency control (OCC) [kungOptimisticMethodsConcurrency1981] established the read-validate-write paradigm; Silo [tuSpeedyTransactionsMulticore2013] demonstrated OCC can achieve nearly 700,000700{,}000 transactions per second on multicore hardware. For high-contention workloads, Doppel [narulaPhaseReconciliationContended2014] introduced phase reconciliation, and deterministic approaches like BOHM [faleiroRethinkingSerializableMultiversion2015] pre-compute dependencies to enable parallel execution. Ira is orthogonal to these approaches: even when transactions execute in parallel, each thread must manage its cache. By providing exact future access information, Ira enables optimal caching within each parallel execution context. Appendix˜F provides extended discussion.
7 Conclusion
Disclaimer
Case studies, comparisons, statistics, research and recommendations are provided “AS IS” and intended for informational purposes only and should not be relied upon for operational, marketing, legal, technical, tax, financial or other advice. Visa Inc. neither makes any warranty or representation as to the completeness or accuracy of the information within this document, nor assumes any liability or responsibility that may result from reliance on such information. The Information contained herein is not intended as investment or legal advice, and readers are encouraged to seek the advice of a competent professional where such advice is required.
These materials and best practice recommendations are provided for informational purposes only and should not be relied upon for marketing, legal, regulatory or other advice. Recommended marketing materials should be independently evaluated in light of your specific business needs and any applicable laws and regulations. Visa is not responsible for your use of the marketing materials, best practice recommendations, or other information, including errors of any kind, contained in this document.
Appendix A Ethical Considerations
Stakeholders. We identify the following stakeholders potentially affected by this research: (1) Ethereum node operators who may deploy Ira-L to accelerate block validation; (2) blockchain users whose transactions are validated by these nodes; (3) block proposers who would generate hints; and (4) the broader Ethereum ecosystem, including validators and protocol developers.
Impacts. This work is a performance optimization for an existing, deployed system. It does not involve human subjects, does not collect personal data, and does not introduce new attack vectors. Our research uses only publicly available blockchain data (Ethereum mainnet blocks) that is accessible to any network participant. The optimization benefits all validators equally and does not create asymmetric advantages that could centralize validation power.
Mitigations. We considered whether hint-based prefetching could enable new attacks:
  • Malicious hints: A dishonest proposer could send incorrect hints. By design, hints are advisory and affect only performance, not correctness; validators independently verify all state transitions regardless of hints.
  • Privacy leakage: Hints reveal which storage keys a block accesses. However, this information is already publicly observable on-chain after block confirmation; hints do not expose information beyond what is already public.
Decision. Given that this research (1) introduces no new attack vectors, (2) uses only public data, (3) provides clear benefits to infrastructure operators, and (4) includes safeguards ensuring hints cannot affect correctness, we determined that publication is appropriate.
Appendix B Discussion
@@
Deployment considerations.
Layer-1 optional. Existing database systems can deploy Ira without enforcing hint correctness as part of the underlying synchronization protocol. Blockchains that implement this need not enforce correctness as part of the block proposal mechanism as the hints are advisory and affect only performance.
@@
Corrupted hints. A peer delivering a corrupted or malicious hint is analogous to a peer delivering a corrupt block: in both cases, the recipient detects the fault and discards the invalid data. For hints, detection is straightforward: decompression failure (Zstandard integrity check), a malformed entry count, or a cache miss during execution (caught by the crash-on-miss invariant, Section˜4.2) all cause the backup to discard the hint and fall back to standard replay. Because hints are advisory and never affect state correctness, the impact of a corrupted hint is bounded to a single block’s replay latency.
We treat the problem of identifying and penalizing peers that deliver corrupt hints as orthogonal to our work. Ethereum’s networking stack already provides mechanisms for this: the consensus layer’s gossipsub peer scoring [SpecsPubsubGossipsub] tracks peer behavior and disconnects peers whose scores fall below a threshold, while the execution layer’s devp2p protocol supports explicit disconnect codes for protocol violations [EthereumDevp2p2026]. These same mechanisms can be extended to penalize peers that consistently deliver invalid hints, requiring no changes to Ira itself.
@@
Appendix C Reth Architecture
@@
Ethereum uses two clients: a consensus client that follows a beacon chain to determine the blockchain that is valid, and an execution client that executes the blocks from the consensus client. They work together to keep a node in sync, validate incoming blocks, and to propose new blocks. This section describes the architecture of Reth, a Rust-based high performance execution client, in particular, we focus on the block validation and execution architecture relevant to understanding our hint-based optimization. Figure˜14 presents a schematic summary of the architecture that we are interested in.
C.1 System components
@@ C.3 MDBX database
C.4 Reth Baseline Architecture
@@ Appendix D Ira-L Architecture
D.1 Reth Primary
reth-primary is the hint generator that simulates a block proposer. It replays blocks, collects all state keys accessed during execution, and writes hints that enable other backups to prefetch state for faster replay. Figure˜15 presents an illustration of our architecture.
Components.
@@ D.1 Reth Primary
D.2 Reth Backup
reth-backup is the hint consumer that simulates a validator receiving hints from the proposer. It prefetches all required state before execution, so that block processing incurs zero disk I/O. Figure˜16 presents an illustration of our architecture.
@@ D.2 Reth Backup
Appendix E Additional Details
Table˜5 provides detailed bucketed statistics for the intra-block access frequency distribution shown in Figure˜4. Table˜6 provides the detailed breakdown for the contract concentration shown in Figure˜5. Table˜7 provides the detailed breakdown for the key lifespan distribution shown in Figure˜6. Table˜8 provides per-block statistics shown in Figure˜7. Section˜E.2 provides the source annotation distribution referenced in Section˜4.1. Section˜E.3 provides the tabulated backup speedup and tail behavior data. Section˜E.4 provides a detailed analysis of hint cost scaling and variance.
@@ Appendix E Additional Details
E.1 Per-Block Operation Distribution
Figure˜17 shows the per-block breakdown of state operations, complementing Table˜2. Storage operations consistently dominate across all blocks.
E.2 Source Annotation Distribution
@@
E.3 Backup Speedup Details
Table˜10 provides the per-block speedup percentiles summarizing Figure˜11. Table˜11 provides the tail behavior breakdown referenced in Section˜5.3.2.
@@ E.3 Backup Speedup Details
E.4 Hint Cost Variance Analysis
[…]
Figure 18: CDF of per-block hint cost in absolute time (ms). Construction time varies widely with block size, while serialization clusters tightly around ${\sim}12.5{}$ ms.
[…]
Table˜12 shows how the hint fraction changes with block working set size. The median hint fraction decreases monotonically from $24.8{}\%$ for small blocks (<<500 keys) to $9.7{}\%$ for large blocks (>>5K keys), even though absolute hint time increases. This occurs because execution time grows faster than hint time for larger blocks: large blocks perform more I/O, and the I/O-dominated execution time scales super-linearly with working set size while hint construction scales linearly (Figure˜19).